Translate

New Live Courses

๐Ÿš€ *Gen AI Engineer Telugu (Production Focused)*
๐Ÿ—“️ *Date:* 30th Sept 2026, 07:00AM IST
๐Ÿ“ *Register Now!:*
https://www.vlrt.in/gr
๐Ÿ‘ฅ *Join WA Community:*
https://www.vlrt.in/gw
๐Ÿ’ก*Course Content:*
https://www.vlrt.in/ai
▶️ *Demo Videos:*
https://www.vlrt.in/gv

๐Ÿš€ *Service now Admin/ Development (ITSM)FREE Demo in Telugu*
๐Ÿ—“️ *Date:* 30th Sept 2026, 08:00AM IST
๐Ÿ“ *Register Now!:*
https://www.vlrt.in/7r
๐Ÿ‘ฅ *Join WA Community:*
https://www.vlrt.in/7w
๐Ÿ’ก*Course Content:*
https://www.vlrt.in/7c
▶️ *Demo Videos:*
https://www.vlrt.in/7v

๐Ÿš€ *Vulnerability Management Training Demo*
๐Ÿ—“️ *Date:* 30th Sept 2026, 09:00 AM IST
๐Ÿ“*Register Now!:*
https://www.vlrt.in/vr
๐Ÿ‘ฅ *Join Community:*
https://www.vlrt.in/cw
๐Ÿ’ก *Course Content:*
https://www.vlrt.in/vc
▶️ *Demo Videos:*
https://www.vlrt.in/Vm

Monday, 26 February 2024

A problem is given to three students whose chances of solving it are 1/2, 1/3, and 1/4 respectively. What is the probability that the problem will be solved?

 Q: A problem is given to three students whose chances of solving it are 1/2, 1/3, and 1/4 respectively. What is the probability that the problem will be solved?

  • A) 7/12

  • B) 1/2

  • C) 3/4

  • D) 1/4

Answer: C) 3/4

Explanation:

Here's one way to solve this problem:

  1. Probability of NOT solving: Find the probability that none of the students can solve the problem.

  • Student 1: (1 - 1/2) = 1/2

  • Student 2: (1 - 1/3) = 2/3

  • Student 3: (1 - 1/4) = 3/4

  1. Probability of NONE solving: Multiply the individual probabilities of not solving:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. Probability of AT LEAST ONE solving: Subtract the probability of none solving from 1 (representing certainty):

  • 1 - 1/4 = 3/4

Therefore, the probability that the problem will be solved is 3/4.


Hindi

เคช्เคฐเคถ्เคจ: เคเค• เคธเคฎเคธ्เคฏा เคคीเคจ เค›ाเคค्เคฐों เค•ो เคฆी เค—เคˆ เคนै เคœिเคจเค•े เคนเคฒ เค•เคฐเคจे เค•ी เคธंเคญाเคตเคจा เค•्เคฐเคฎเคถः 1/2, 1/3 เค”เคฐ 1/4 เคนै। เค•्เคฏा เคธंเคญाเคตเคจा เคนै เค•ि เคธเคฎเคธ्เคฏा เคนเคฒ เคนो เคœाเคฏेเค—ी?

  • เค) 7/12

  • เคฌी) 1/2

  • เคธी) 3/4

  • เคกी) 1/4

เค‰เคค्เคคเคฐ: เคธी) 3/4

เคธ्เคชเคท्เคŸीเค•เคฐเคฃ:

เค‡เคธ เคธเคฎเคธ्เคฏा เค•ो เคนเคฒ เค•เคฐเคจे เค•ा เคเค• เคคเคฐीเค•ा เคฏเคนां เคฆिเคฏा เค—เคฏा เคนै:

  1. เคนเคฒ เคจ เค•เคฐเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा: เคช्เคฐाเคฏिเค•เคคा เคœ्เคžाเคค เค•ीเคœिเค เค•ि เค•ोเคˆ เคญी เค›ाเคค्เคฐ เคธเคฎเคธ्เคฏा เค•ा เคธเคฎाเคงाเคจ เคจเคนीं เค•เคฐ เคธเค•เคคा।

  • เคตिเคฆ्เคฏाเคฐ्เคฅी 1: (1 - 1/2) = 1/2

  • เค›ाเคค्เคฐ 2: (1 - 1/3) = 2/3

  • เค›ाเคค्เคฐ 3: (1 - 1/4) = 3/4

  1. เค•िเคธी เค•े เคญी เคนเคฒ เคจ เค•เคฐเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा: เคนเคฒ เคจ เค•เคฐ เคชाเคจे เค•ी เคต्เคฏเค•्เคคिเค—เคค เคธंเคญाเคตเคจाเค“ं เค•ो เค—ुเคฃा เค•เคฐें:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. เค•เคฎ เคธे เค•เคฎ เคเค• เคฌाเคฐ เคนเคฒ เค•เคฐเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा: เค•ोเคˆ เคญी เคนเคฒ เคจ เค•เคฐ เคชाเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा เค•ो 1 เคธे เค˜เคŸाเคँ (เคจिเคถ्เคšिเคคเคคा เค•ा เคช्เคฐเคคिเคจिเคงिเคค्เคต เค•เคฐเคคे เคนुเค):

  • 1 - 1/4 = 3/4

เค…เคคः เคธเคฎเคธ्เคฏा เค•े เคนเคฒ เคนो เคœाเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा 3/4 เคนै।


Telugu

 เฐช్เฐฐ: เฐฎుเฐ—్เฐ—ుเฐฐు เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅుเฐฒเฐ•ు เฐ’เฐ• เฐธเฐฎเฐธ్เฐฏ เฐ‡เฐต్เฐตเฐฌเฐกింเฐฆి, เฐฆాเฐจి เฐชเฐฐిเฐท్เฐ•ాเฐฐ เฐ…เฐตเฐ•ాเฐถాเฐฒు เฐตเฐฐుเฐธเฐ—ా 1/2, 1/3 เฐฎเฐฐిเฐฏు 1/4. เฐธเฐฎเฐธ్เฐฏ เฐชเฐฐిเฐท్เฐ•ాเฐฐเฐฎเฐฏ్เฐฏే เฐธంเฐญాเฐต్เฐฏเฐค เฐเฐฎిเฐŸి?

  • เฐŽ) 7/12

  • เฐฌి) 1/2

  • เฐธి) 3/4

  • เฐกి) 1/4

เฐธเฐฎాเฐงాเฐจం: เฐธి) 3/4

เฐตిเฐตเฐฐเฐฃ:

เฐˆ เฐธเฐฎเฐธ్เฐฏเฐจు เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐกాเฐจిเฐ•ి เฐ‡เฐ•్เฐ•เฐก เฐ’เฐ• เฐฎాเฐฐ్เฐ—ం เฐ‰ంเฐฆి:

  1. เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐจి เฐธంเฐญాเฐต్เฐฏเฐค: เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅుเฐฒెเฐตเฐฐూ เฐธเฐฎเฐธ్เฐฏเฐจు เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐฒేเฐจి เฐธంเฐญాเฐต్เฐฏเฐคเฐจు เฐ•เฐจుเฐ—ొเฐจంเฐกి.

  • เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅి 1: (1 - 1/2) = 1/2

  • เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅి 2: (1 - 1/3) = 2/3

  • เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅి 3: (1 - 1/4) = 3/4

  1. เฐเฐฆీ เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐจి เฐธంเฐญాเฐต్เฐฏเฐค: เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐจి เฐต్เฐฏเฐ•్เฐคిเฐ—เฐค เฐธంเฐญాเฐต్เฐฏเฐคเฐฒเฐจు เฐ—ుเฐฃింเฐšంเฐกి:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. เฐ•เฐจీเฐธం เฐ’เฐ• เฐชเฐฐిเฐท్เฐ•ాเฐฐం เฐฏొเฐ•్เฐ• เฐธంเฐญాเฐต్เฐฏเฐค: เฐเฐฆీ เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐจి เฐธంเฐญాเฐต్เฐฏเฐคเฐจు 1 เฐจుంเฐกి เฐคీเฐธిเฐตేเฐฏంเฐกి (เฐจిเฐถ్เฐšเฐฏเฐคเฐจు เฐธూเฐšిเฐธ్เฐคుంเฐฆి):

  • 1 - 1/4 = 3/4

เฐ…ంเฐฆుเฐตเฐฒ్เฐฒ, เฐธเฐฎเฐธ్เฐฏ เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐฌเฐกే เฐธంเฐญాเฐต్เฐฏเฐค 3/4.


Tamil

 เฎ•ே: เฎ’เฎฐு เฎšிเฎ•்เฎ•เฎฒைเฎค் เฎคீเฎฐ்เฎช்เฎชเฎคเฎฑ்เฎ•ாเฎฉ เฎตாเฎฏ்เฎช்เฎชு เฎฎுเฎฑைเฎฏே 1/2, 1/3 เฎฎเฎฑ்เฎฑுเฎฎ் 1/4 เฎ†เฎ•ிเฎฏ เฎฎூเฎฉ்เฎฑு เฎฎாเฎฃเฎตเฎฐ்เฎ•เฎณுเฎ•்เฎ•ு เฎตเฎดเฎ™்เฎ•เฎช்เฎชเฎŸுเฎ•ிเฎฑเฎคு. เฎชிเฎฐเฎš்เฎšเฎฉை เฎคீเฎฐுเฎฎ் เฎจிเฎ•เฎด்เฎคเฎ•เฎตு เฎŽเฎฉ்เฎฉ?

  • A) 7/12

  • B) 1/2

  • C) 3/4

  • D) 1/4

เฎชเฎคிเฎฒ்: C) 3/4

เฎตிเฎณเฎ•்เฎ•เฎฎ்:

เฎ‡เฎจ்เฎค เฎšிเฎ•்เฎ•เฎฒை เฎคீเฎฐ்เฎ•்เฎ• เฎ’เฎฐு เฎตเฎดி เฎ‡เฎ™்เฎ•ே:

  1. เฎคீเฎฐ்เฎ•்เฎ•เฎช்เฎชเฎŸாเฎฎเฎฒ் เฎ‡เฎฐுเฎช்เฎชเฎคเฎฑ்เฎ•ாเฎฉ เฎจிเฎ•เฎด்เฎคเฎ•เฎตு: เฎฎாเฎฃเฎตเฎฐ்เฎ•เฎณ் เฎฏாเฎฐுเฎฎ் เฎšிเฎ•்เฎ•เฎฒைเฎค் เฎคீเฎฐ்เฎ•்เฎ• เฎฎுเฎŸிเฎฏாเฎค เฎจிเฎ•เฎด்เฎคเฎ•เฎตைเฎ•் เฎ•เฎฃ்เฎŸเฎฑிเฎฏเฎตுเฎฎ்.

  • เฎฎாเฎฃเฎตเฎฐ் 1: (1 - 1/2) = 1/2

  • เฎฎாเฎฃเฎตเฎฐ் 2: (1 - 1/3) = 2/3

  • เฎฎாเฎฃเฎตเฎฐ் 3: (1 - 1/4) = 3/4

  1. เฎคீเฎฐ்เฎ•்เฎ•เฎช்เฎชเฎŸாเฎค เฎจிเฎ•เฎด்เฎคเฎ•เฎตு: เฎคீเฎฐ்เฎ•்เฎ•ாเฎค เฎคเฎฉிเฎช்เฎชเฎŸ்เฎŸ เฎจிเฎ•เฎด்เฎคเฎ•เฎตுเฎ•เฎณை เฎชெเฎฐுเฎ•்เฎ•เฎตுเฎฎ்:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. เฎ•ுเฎฑைเฎจ்เฎคเฎชเฎŸ்เฎšเฎฎ் เฎ’เฎฐு เฎคீเฎฐ்เฎตுเฎ•்เฎ•ாเฎฉ เฎจிเฎ•เฎด்เฎคเฎ•เฎตு: เฎŽเฎคுเฎตுเฎฎ் เฎคீเฎฐ்เฎ•்เฎ•เฎช்เฎชเฎŸாเฎฎเฎฒ் เฎ‡เฎฐுเฎช்เฎชเฎคเฎฑ்เฎ•ாเฎฉ เฎจிเฎ•เฎด்เฎคเฎ•เฎตை 1 เฎ‡เฎฒிเฎฐுเฎจ்เฎคு เฎ•เฎดிเฎ•்เฎ•เฎตுเฎฎ் (เฎ‰เฎฑுเฎคிเฎฏாเฎฉ เฎคเฎฉ்เฎฎைเฎฏைเฎ•் เฎ•ுเฎฑிเฎ•்เฎ•ுเฎฎ்):

  • 1 - 1/4 = 3/4

เฎŽเฎฉเฎตே, เฎšிเฎ•்เฎ•เฎฒ் เฎคீเฎฐ்เฎ•்เฎ•เฎช்เฎชเฎŸுเฎตเฎคเฎฑ்เฎ•ாเฎฉ เฎจிเฎ•เฎด்เฎคเฎ•เฎตு 3/4 เฎ†เฎ•ுเฎฎ்.


Spanish

 P: Se les da un problema a tres estudiantes cuyas posibilidades de resolverlo son 1/2, 1/3 y 1/4 respectivamente. ¿Cuรกl es la probabilidad de que se resuelva el problema?

  • A) 7/12

  • b) 1/2

  • c) 3/4

  • D) 1/4

Respuesta: C) 3/4

Explicaciรณn:

Aquรญ hay una forma de resolver este problema:

  1. Probabilidad de NO resolver: Encuentre la probabilidad de que ninguno de los estudiantes pueda resolver el problema.

  • Estudiante 1: (1 - 1/2) = 1/2

  • Estudiante 2: (1 - 1/3) = 2/3

  • Estudiante 3: (1 - 1/4) = 3/4

  1. Probabilidad de que NINGUNO resuelva: Multiplique las probabilidades individuales de no resolver:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. Probabilidad de que AL MENOS UNO resuelva: Reste la probabilidad de que ninguno resuelva de 1 (que representa certeza):

  • 1 - 1/4 = 3/4

Por tanto, la probabilidad de que el problema se resuelva es 3/4.


French

 Q : Un problรจme est confiรฉ ร  trois รฉlรจves dont les chances de le rรฉsoudre sont respectivement de 1/2, 1/3 et 1/4. Quelle est la probabilitรฉ que le problรจme soit rรฉsolu ?

  • A) 7/12

  • B) 1/2

  • C) 3/4

  • D) 1/4

Rรฉponse : C) 3/4

Explication:

Voici une faรงon de rรฉsoudre ce problรจme :

  1. Probabilitรฉ de NE PAS rรฉsoudre : Trouvez la probabilitรฉ qu’aucun des รฉlรจves ne puisse rรฉsoudre le problรจme.

  • ร‰tudiant 1 : (1 - 1/2) = 1/2

  • ร‰tudiant 2 : (1 - 1/3) = 2/3

  • ร‰lรจve 3 : (1 - 1/4) = 3/4

  1. Probabilitรฉ de ne pas rรฉsoudre : Multipliez les probabilitรฉs individuelles de ne pas rรฉsoudre :

  • (1/2) * (2/3) * (3/4) = 1/4

  1. Probabilitรฉ d’AU MOINS UNE rรฉsolution : soustrayez la probabilitรฉ qu’aucune solution ne soit rรฉsolue de 1 (reprรฉsentant la certitude) :

  • 1 - 1/4 = 3/4

La probabilitรฉ que le problรจme soit rรฉsolu est donc de 3/4.


No comments:

Post a Comment

Note: only a member of this blog may post a comment.